Rose purchased a total of 17 books and toys for the Rupert Play School. Each book costs $11 and each toy costs $6. How many books and toys did she buy for $157?

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

book=x
toy=17-x, since they both add to 17
x*11+(17-x)*6=157
11x+102-6x=157
5x=55
x=11 books ($121)
17-x=6 toys ($36)

Answer by greenestamps(13206)



Using logical reasoning instead of formal algebra....

(1) If all 17 items were toys, the cost would be 17*$6 = $102.
(2) The actual total cost was $55 more than that.
(3) Each book costs $5 more than each toy.
(4) Therefore, the number of books was 55/5 = 11.
(5) That of course means 6 books.

CHECK: 11($11)+6($6) = $121+$36 = $157